pandasurvive:

The length of a rectangular yard is 4 ft longer than its width w. The area of the yard is 525 ft^2. What are the dimensions of the yard?

Ultrilliam:

heres a similar question that should help you solve this problem: questioncove.com/updates/54d530cbe4b06e8258bdb065 if that doesn't help you may want to tag hero ^_^

pandasurvive:

@Ultrilliam @Hero I put this but my teacher marked me wrong

1 attachment
Hero:

@pandasurvive, what is the question asking?

pandasurvive:

Just for the dimensions of the yard.

Hero:

Meaning what exactly?

Hero:

How does one break down the description for "dimensions of a yard" in this case?

pandasurvive:

It doesnt say....

Hero:

"What are the dimensions of the yard?" breaks down to two questions: 1. What is the length of the yard? 2. What is the width of the yard?

Hero:

You figured out that we have to use the formula A = lw to answer the question. You also figured out A = 525 l = w + 4 w = w You even solved for w: 525 = (w + 4)(w) 525 = w^2 + 4w 0 = w^2 + 4w - 525 0 = (w - 21)(w + 25) 0 = w - 21 w = 21 In other words you found the width, but you forgot to find the length. Also, for every word problem, your final answer must be expressed as a complete sentence.

Hero:

@pandasurvive

pandasurvive:

I never knew that last part.

Hero:

Yes, and never forget it.

pandasurvive:

L=-25?

Hero:

What does "The length of the yard is 4 ft longer than the width" mean to you?

pandasurvive:

That its not negative

pandasurvive:

25

Hero:

Yes correct.

Hero:

The width of the yard is 21 ft. The length of the yard (4 more than that) must be +25 ft.

pandasurvive:

Thanks

Hero:

You're most welcome.